# Brain network science modelling of sparse neural networks enables Transformers and LLMs to perform as fully connected

### v2

@@ Appendix I Integral of Sigmoid Density Decay and Cubic Density Decay
In this section, we show the formula for the proposed Sigmoid Density Decay and the Cubic decay implemented in GraNet [42] and GMP [55].
@@
[…]
@@
[…]
@@
Inspired by GMP [55, 43], we incorporate historical weights into our CHTs and CHTss implementation. During training, we maintain a historical weight matrix that records previously learned weights throughout the training process. When CHTs and CHTss predict new links, we initialize them using their corresponding historical weights - specifically, the values they held before being pruned. In this way, CHTs and CHTss enable weight recovery with preserved memory, allowing the model to retain valuable prior information.
@@
1:  Binary bipartite adjacency matrix $A_{UV}\!\in\!\{0,1\}^{m\times n}$triangleright $U$–to–$V$ edges
@@
4:       $A_{VU}\leftarrow A_{UV}^{\top}$triangleright $V\!\to\!U$ edges
5:       $d_{U}\leftarrow$ row-sum$(A_{UV})$
6:       $d_{V}\leftarrow$ row-sum$(A_{VU})$
@@
8:       $UU\leftarrow A_{UV}A_{VU}$triangleright $U\!\to\!V\!\to\!U$
9:       $VV\leftarrow A_{VU}A_{UV}$triangleright $V\!\to\!U\!\to\!V$
@@
15:                $e_{UU}[i,j]\leftarrow\dfrac{UU[i,j]+1}{ext+1}$triangleright # According to Equation (3)
@@
21:                $e_{VV}[a,b]\leftarrow\dfrac{VV[a,b]+1}{ext+1}$triangleright # According to Equation (3)
@@
We present a comparison of CHTs, CHTss, and fully connected network on language modeling tasks using the LLaMA-1B model on Table 14. The results clearly demonstrate that CHTs consistently outperform the fully connected (FC) baseline at 70%, even at a high sparsity of 95%, CHTss achieves a perplexity of 16.51, which is remarkably close to the FC baseline.
@@
We also compare the performance of CHTs, CHTss, and the fully connected network on zero-shot tasks, as summarized in Table 15. The results demonstrate that CHTs generally outperforms the fully connected baseline at 90% and 95% sparsity levels. This suggests that CHTs enhances the model’s generalization capability on previously unseen data.
Table 15: Zero-shot evaluation of LLaMA-1B across GLUE and SuperGLUE. ACC scores are shown. Values are mean ± sd over 5 seeds (lm-eval). Red values indicate the top scores over models and sparsities for a benchmark.
@@
Appendix M Experiments compute resources

### v3

@@ Appendix I Integral of Sigmoid Density Decay and Cubic Density Decay
In this section, we show the formula for the proposed Sigmoid Density Decay and the Cubic decay implemented in GraNet [liu2021granet] and GMP [han2015gmp].
@@
$\begin{split}\int_{t_{0}}^{t_{f}}(s_{i}-s_{f})\left(1-\frac{t-t_{0}}{n\Delta t}\right)^{3}dt\\ =\frac{1}{4}(s_{i}-s_{f})(t_{f}-t_{0}).\end{split}$ (10)
@@
$\begin{split}\int_{t_{0}}^{t^{\prime}_{f}}(s_{i}-s_{f})\left(\frac{1}{1+e^{-k\left(t-\frac{t^{\prime}_{f}+t_{0}}{2}\right)}}\right)dt\\ =\frac{(s_{i}-s_{f})(t^{\prime}_{f}-t_{0})}{2}.\end{split}$ (11)
@@
Inspired by GMP [han2015gmp, zhu2017gmp], we incorporate historical weights into our CHTs and CHTss implementation. During training, we maintain a historical weight matrix that records previously learned weights throughout the training process. When CHTs and CHTss predict new links, we initialize them using their corresponding historical weights - specifically, the values they held before being pruned. In this way, CHTs and CHTss enable weight recovery with preserved memory, allowing the model to retain valuable prior information.
@@
1:  Binary bipartite adjacency matrix $A_{UV}\!\in\!\{0,1\}^{m\times n}$ $\triangleright$ UU–to–$V$ edges
@@
4:    $A_{VU}\leftarrow A_{UV}^{\top}$ $\triangleright$ $V\!\to\!U$ edges
[…]
@@
8:    $UU\leftarrow A_{UV}A_{VU}$ $\triangleright$ $U\!\to\!V\!\to\!U$
9:    $VV\leftarrow A_{VU}A_{UV}$ $\triangleright$ $V\!\to\!U\!\to\!V$
@@
15:       $e_{UU}[i,j]\leftarrow\dfrac{UU[i,j]+1}{ext+1}$ $\triangleright$ # According to Equation (3)
@@
21:       $e_{VV}[a,b]\leftarrow\dfrac{VV[a,b]+1}{ext+1}$ $\triangleright$ # According to Equation (3)
@@
We present a comparison of CHTs, CHTss, and fully connected network on language modeling tasks using the LLaMA-1B model on Table 17. The results clearly demonstrate that CHTs outperform the fully connected (FC) baseline at 70%, even at a high sparsity of 95%, CHTss achieves a perplexity of 16.51, which is remarkably close to the FC baseline.
@@
We compare CHTs, CHTss, and a fully connected (FC) network on zero-shot benchmarks, as summarized in Table 20. At 90% and 95% sparsity, CHTs often surpasses the FC baseline, suggesting improved generalization to unseen data. At 70% sparsity, CHTs is comparable to FC, whereas at higher sparsity (90–95%) it can exceed FC in accuracy. This pattern appears at odds with perplexity trends, so we provide additional analysis for readers evaluating pretrained models on GLUE/SuperGLUE in a zero-shot setting.
A key confound is label imbalance in several tasks: some datasets contain markedly skewed positive/negative distributions. In practice, we observed that models sometimes emit nearly the same label for most inputs, yielding deceptively strong accuracy if that label matches the dataset majority class. Table 18 reports the class ratios for representative tasks.
Table 18: Label imbalance in selected GLUE/SuperGLUE tasks (positive vs. negative counts and proportions).
Dataset Pos (count, %) Neg (count, %)
CoLA 721 (69.13%) 322 (30.87%)
MRPC 279 (68.38%) 129 (31.62%)
QQP 14,885 (36.82%) 25,545 (63.18%)
BoolQ 2,033 (62.17%) 1,237 (37.83%)
To mitigate this issue, we report Matthews correlation coefficient (MCC) instead of accuracy. MCC is robust to class imbalance and better reflects the quality of binary predictions. Table 19 shows averaged MCC over GLUE/SuperGLUE under multiple sparsity regimes.
Table 19: Average MCC on zero-shot GLUE/SuperGLUE. FC denotes the dense fully connected model.
Metric FC 70% CHTs 70% CHTss 90% CHTs 90% CHTss 95% CHTs    95% CHTss
AVG MCC 0.031 0.028 0.033 0.031 0.022 0.003-0.003    0.007-0.007
The 70% sparse models (MCC $0.028$ for CHTs; $0.033$ for CHTss) are on par with the dense FC baseline (MCC $0.031$), but MCC declines as sparsity increases; at 95% sparsity, negative MCC indicates predictions are anticorrelated with ground truth (worse than random). We attribute part of this degradation to compute-limited pretraining budgets (e.g., up to 8.9B tokens for LLaMA-1B in our runs), which are insufficient to reach the regime where zero-shot performance is reliable at extreme sparsity. Recommendation: for zero-shot evaluation on imbalanced datasets, prefer MCC (or similarly balanced metrics) over raw accuracy, especially when comparing across sparsity levels and training budgets.
Table 20: Zero-shot evaluation of LLaMA-1B across GLUE and SuperGLUE. ACC scores are shown. Values are mean ± sd over 5 seeds (lm-eval [eval-harness]). Red values indicate the top scores over models and sparsities for a benchmark.
@@
Appendix M Extra Related Works
More recently, BiDST has elegantly reframed DST as a bi-level optimization problem to co-optimize the weights and the sparsity mask simultaneously [pmlr-v235-ji24a]. Other methods improve the training dynamics itself, such as AC/DC, which alternates between sparse and dense phases [peste2021acdc], or Powerpropagation, which introduces a sparsity-inducing weight reparameterization [schwarz2021powerpropagation]. To enhance generalization in the often-chaotic loss landscape of sparse models, S2-SAM provides an efficient, plug-and-play sharpness-aware optimizer [ji2024s2sam]. Another thrust of research adapts DST for specific architectural or hardware advantages. For example, SLaK utilizes dynamic sparsity to enable massive, high-performance kernels in CNNs [liu2022slak]. Chase provides a practical method to translate unstructured dynamic sparsity into hardware-friendly, channel-level sparsity [yin2023chase]. Addressing network health, NeurRev proposes a mechanism to revitalize “dormant neurons” that can emerge during training [li2024neurrev].
Appendix N Experiments compute resources